OpenStudy (theviper):

Plz help:) \[\Huge{\color{red}{\mathbb{a^2+b^2=?}}}\]\[\LARGE{\color{green}{\text{a+b=7 and ab=12}}}\]

OpenStudy (btaylor):

\[(a+b)^2 = a^2 + 2ab + b^2\]So \[a^2 + b^2 = 49-2(12) = 49-24=25\]

OpenStudy (theviper):

\[\Huge{\color{gold}{\star \star \star}}\]Yeah! that's the ans & I got it!